# NuLite - Lightweight and Fast Model for Nuclei Instance Segmentation and Classification

### v1

@@ 4.1 Datasets
The PanNuke dataset is the primary resource for training and evaluating our model. It comprises 189,744 annotated nuclei across 7,904 images, each of size 256$\times$256 pixels, spanning 19 distinct tissue types and categorized into five unique cell classes. These cell images were captured at a 40$\times$ magnification with a fine resolution of 0.25 $\mu$m/px. Notably, the dataset exhibits a significant class imbalance; particularly, the nuclei class of dead cells is markedly underrepresented, evident from the nuclei and tissue class statistics.
@@
The MoNuSeg dataset is employed as a supplementary resource for nuclei segmentation. Unlike PanNuke, MoNuSeg is considerably smaller and does not categorize nuclei into various classes. In this study, only the test subset of MoNuSeg is used to assess our model. This subset includes 14 high-resolution images (1000 $\times$ 1000 px) captured at 40$\times$ magnification and a resolution of 0.25 $\mu$m/px, containing over 7,000 annotated nuclei spanning seven organ types (kidney, lung, colon, breast, bladder, prostate, and brain) across various disease states. Due to the absence of nuclei labels, classification performance cannot be evaluated with this dataset.
@@
The CoNSeP dataset, curated by Graham et al., comprises 41 H&E-stained colorectal adenocarcinoma whole slide images (WSIs) at a resolution of 0.25 $\mu$m/px, resized to 1024 $\times$ 1024 px to facilitate processing. This diverse dataset features stromal, epithelial, muscular, collagen, adipose, and tumorous regions. It also includes a variety of nuclei types derived from different originating cells, such as normal epithelial, dysplastic epithelial, inflammatory, necrotic, muscular, fibroblast, and miscellaneous nuclei types, including necrotic and mitotic cells, which aids in comprehensive phenotypic analysis.
@@
The GLySAC dataset, short for Gastric Lymphocyte Segmentation and Classification, focuses on segmenting and classifying nuclei within gastric pathology. It contains 59 H&E stained image tiles, each 1000x1000 pixels, sourced from gastric adenocarcinoma WSIs and captured at a 40$\times$ magnification using an Aperio digital scanner. The dataset encapsulates a total of 30,875 nuclei, categorized into three primary groups: Lymphocytes (12,081 nuclei), Epithelial nuclei (12,287 nuclei), encompassing both cancerous and normal cells, and Miscellaneous other nuclei types (6,507 nuclei).
@@ 4.2 Evaluation metrics
SQ assesses the accuracy of the segmentation for the detected nuclei, computed as the mean IoU (Intersection over Union) of matched segments, as denoted in Equation 7, where $y$ and $\hat{y}$ denote the ground truth and predicted segments, respectively.
@@
Therefore, panoptic quality is definite as the product of detection quality and segmentation quality as denoted in equation 8.
@@
Several metrics commonly utilized in machine learning were employed to evaluate our model’s classification performance. Precision ($P$), which quantifies the accuracy of the model’s positive predictions, is defined in the equation environment as follows:
@@
where $TP$ represents true positives and $FP$ stands for false positives. Recall ($R$), also known as sensitivity, measures the ability of the model to detect all relevant instances encapsulated by:
@@
with $FN$ indicating false negatives. The F1 Score, a harmonic mean of precision and recall that provides a balance between these metrics, is crucial in the presence of uneven class distribution and is expressed by:
@@
Accuracy, indicating the overall correctness of the model, is formulated as:
@@
To accommodate detailed performance assessment in multi-class settings, the F1 Score is further refined through equations that include terms for each class $c$ and $d$, illustrating both traditional components and inter-class effects:
$F1_{c}=\frac{2(T_{Pc}+T_{Nc})}{2(T_{Pc}+T_{Nc})+2FP_{c}+2FN_{c}+FP_{d}+FN_{d}}$ (13)
@@
These refinements enable a more comprehensive evaluation of the model discriminatory capacity between classes, essential in contexts where class imbalance or differential inter-class errors significantly impact performance.
@@ 4.3 Results on PanNuke
DIST 0.439 0.29 0.343 0.275 0.000
Mask-RCNN 0.472 0.403 0.290 0.300 0.069
Micro-Net 0.504 0.442 0.333 0.334 0.051
HoVer-Net 0.551 0.491 0.417 0.388 0.139
HoVer-UNet 0.524 0.478 0.401 0.379 0.076
CellViT256 0.567 0.559 0.405 0.405 0.144
CellViT-SAM-H 0.581 0.583 0.417 0.423 0.149
NuLite-T 0.567 0.562 0.418 0.401 0.155
@@
Tissue HoVer-Net CellViT256 CellViT-SAM-H NuLite-T12 NuLite-SA24 NuLite-SA36
@@
No Reparameterized FastViT-T8 0.6520 0.4808 0.8211 0.5668 0.5512 0.4196 0.3967 0.1479 0.6902 0.7033 0.5788 0.5164 0.3719
FastViT-T12 0.6546 0.4825 0.8225 0.5666 0.5612 0.4183 0.4009 0.1536 0.6910 0.7084 0.5764 0.5173 0.3756
FastViT-S12 0.6550 0.4862 0.8227 0.5719 0.5600 0.4241 0.4055 0.1354 0.6944 0.7073 0.5806 0.5176 0.3585
FastViT-SA12 0.6576 0.4874 0.8228 0.5732 0.5575 0.4243 0.4065 0.1417 0.6914 0.7040 0.5837 0.5218 0.3604
FastViT-SA24 0.6610 0.4893 0.8246 0.5773 0.5597 0.4215 0.4087 0.1504 0.6989 0.7193 0.5798 0.5216 0.3627
FastViT-SA36 0.6620 0.4900 0.8246 0.5764 0.5661 0.4147 0.4123 0.1526 0.7010 0.7183 0.5793 0.5247 0.3657
FastViT-MA36 0.6621 0.4912 0.8254 0.5779 0.5732 0.4199 0.4119 0.1442 0.7005 0.7249 0.5809 0.5261 0.3540
Reparameterized FastViT-T8 0.6520 0.4810 0.8215 0.5671 0.5514 0.4194 0.3970 0.1470 0.6904 0.7035 0.5782 0.5165 0.3719
FastViT-T12 0.6548 0.4827 0.8225 0.5667 0.5615 0.4183 0.4011 0.1545 0.6911 0.7084 0.5768 0.5176 0.3760
FastViT-S12 0.6551 0.4864 0.8228 0.5718 0.5606 0.4244 0.4056 0.1361 0.6944 0.7075 0.5809 0.5177 0.3593
FastViT-SA12 0.6575 0.4873 0.8227 0.5732 0.5567 0.4241 0.4064 0.1406 0.6913 0.7037 0.5840 0.5221 0.3606
FastViT-SA24 0.6611 0.4895 0.8245 0.5772 0.5606 0.4217 0.4086 0.1505 0.6990 0.7192 0.5791 0.5216 0.3629
FastViT-SA36 0.6621 0.4901 0.8247 0.5766 0.5658 0.4148 0.4123 0.1529 0.7009 0.7184 0.5796 0.5244 0.3664
FastViT-MA36 0.6622 0.4913 0.8253 0.5781 0.5732 0.4201 0.4117 0.1429 0.7006 0.7248 0.5810 0.5263 0.3533
@@
We used the PanNuke dataset to evaluate the performance of our models on nuclei instance segmentation and classification of five distinct cell types: neoplastic, epithelial, inflammatory, connective, and dead cells. We consider the PQ and F-score for each nuclei type and the F-score for detection to perform a robust analysis. Furthermore, we compare our results with DIST, MASK-RCNN, MICRO-Net, HoVer-UNet, CellViT256, and CellViT-SAM-H. In the following results, we consider three NuLite versions: NuLite-T, NuLite-M, and NuLite-H, which respectively use FastViT-T12, FastViT-SA12, and FastViT-SA36 as encoders and, in inference, they are reparameterized as described in [12]. In the ablation study section, we justify the selected encoders and reparameterization. To compare our model with others, we take into account three aspects. First, we analyze the PQ for each nucleus type, reported in 2, then we analyze the F1-score, reported in Table 3. Lastly, we analyze the results regarding binary Panoptic Quality (bPQ) and multi-class Panoptic Quality (mPQ) over each tissue. All results are an average of over three training sessions, as the authors’ dataset suggested. PQ results show that our model is almost the second best and, in the case of inflammatory nuclei, is the best. Furthermore, the results for the different versions of NuLite are pretty similar. Again, analyzing the results in terms of the F1-score, reported in Table 3, NuLite achieves results practically equal to CellViT and outperforms other methods. Lastly, we compare the results over tissues, reported in Table 4, in terms of binary Panoptic Quality (bPQ) and multi-class Panoptic Quality (mPQ). These results prove that our NuLite is practically equal to CellViT-SAM-H, the best model, but more complex and slow than ours, as demonstrated in the following. Moreover, our models outperform CellViT256, the fastest version in the SOTA. In the model complexity analysis paragraph, we discuss the efficiency of NuLite in terms of parameters and the number of multiple additions, proving that it is the fastest and lightest SOTA. Lastly, we also compared our versions with CellViT over tissue types, as shown in Table 4. Again, these results demonstrate that NuLite in all its versions is similar to CellViT in binary panoptic quality (bPQ) and multi-class panoptic quality (mPQ). Therefore, these results show that using NuFast-ViT-H, we obtain the same results of CellViT-SAM-H in terms of PQ and F1-score, except for Epthitelial nuclei. Also, considering a tiny version of NuLite, the results are not different from CellViT-SAM-H and are better or equal to CellViT256, which is the lightest version of it.
@@ 4.4 Ablation study
In this section, we describe the methodology used for selecting our models. We evaluated our NuLite using all versions of FastViT, considering both reparameterized and non-reparameterized variants during the inference phase. The results, as shown in Table 5, indicate that reparameterization does not significantly defect performance. Consequently, we opted for the reparameterized versions due to their superior computational efficiency, as demonstrated in Table 6. To provide a comprehensive representation, we selected three versions of NuLite, namely NuLite-T, NuLite-M, and NuLite-H, corresponding to tiny, medium, and large configurations, respectively. This selection was informed by the evaluation results on the PanNuke dataset, detailed in Table 5. The metrics used for evaluation included binary panoptic quality (PQ), multi-class panoptic quality (mPQ), F1-score, panoptic quality, and F1-score for each nucleus type in the PanNuke dataset. The nucleus types are denoted as Neoplastic (N), Epithelial (E), Inflammatory (I), Connective/Soft Tissue (C), and Dead (D). Upon analyzing the results for the reparameterized models, it is apparent that the performance metrics are closely aligned across most variants. Thus, model selection also considered inference time and model complexity, as detailed in Table 6. First, we exclude the results with FastViT-T8 and FastViT-MA36; the first because despite being the lightest, the results are the worst; the second, because despite being the heaviest, the results are not enough better than the second heaviest model. Then, we chose FastViT-T12 as the backbone for NuLite-T due to its minimal parameter count, FastViT-SA24 as the backbone of NuList-M because it offers a balanced number of parameters and computational complexity without a substantial performance trade-off compared to larger models, and FastViT-SA36 for NuLite-H.
Table 6: Comparison between CellViT and NuLite (ours) on a batch of 4 patches over two input shapes (256, 1024), in terms of the number of parameters, number of multiplications and additions, and estimated size GPU latency
Model Encoder # Parameters shape Multiply-Adds (G) Estimated Total Size (MB) Latency (ms)
CellViT ViT-256 46,750,349 256 132.8913 1,859.98 $72\pm 0.39$
1024 2,125.9402 26,953.06 $1489.1\pm 64.97$
SAM-B 146,094,557 256 200.1459 2,578.66 $95.54\pm 0.2$
1024 3,200.6049 31,184.34 $1437.2\pm 3.48$
SAM-L 367,847,645 256 207.0680 4,050.09 $150.64\pm 0.26$
1024 3,306.5378 37,980.70 $1940.66\pm 1.69$
SAM-H 699,741,149 256 214.2040 6,002.34 $208.68\pm 0.18$
1024 3,413.4114 45,612.96 $2499.22\pm 2.1$
NuLite fastvit_t8 9,237,389 256 14.7031 603.84 $33.34\pm 1.13$
1024 235.2490 9,108.12 $295.56\pm 1$
fastvit_t12 17,157,469 256 26.1658 834.00 $36.96\pm 0.24$
1024 418.6516 12,315.90 $349.84\pm 0.97$
fastvit_s12 19,077,341 256 26.5697 851.64 $35.7\pm 1.23$
1024 425.1140 12,482.97 $353.52\pm 1.09$
fastvit_sa12 21,184,733 256 26.5727 860.87 $35.96\pm 1.28$
1024 425.1318 12,503.99 $357.3\pm 0.91$
fastvit_sa24 31,157,981 256 28.2645 1,020.77 $49.94\pm 1.37$
1024 452.1683 14,464.82 $418.86\pm 1.49$
fastvit_sa36 41,131,229 256 29.9562 1,180.67 $66.32\pm 1.23$
1024 479.2048 16,425.65 $480.48\pm 0.66$
fastvit_ma36 57,837,473 256 42.0835 1,436.89 $74.64\pm 3.08$
1024 673.2023 19,523.35 $742.92\pm 2.12$
NuLite-Rep fastvit_t8 9,210,269 256 14.7005 564.85 $25.8\pm 0.6$
1024 235.2077 8,485.20 $278.72\pm 1.23$
fastvit_t12 17,116,701 256 26.1623 777.81 $26.84\pm 0.79$
1024 418.5965 11,418.23 $324.06\pm 1.09$
fastvit_s12 19,036,573 256 26.5662 795.45 $26.82\pm 0.73$
1024 425.0589 11,585.29 $328.64\pm 0.63$
fastvit_sa12 21,153,181 256 26.5693 806.80 $27.98\pm 0.9$
1024 425.0767 11,639.90 $332.12\pm 0.87$
fastvit_sa24 31,109,149 256 28.2610 928.91 $35.46\pm 1.33$
1024 452.1132 12,996.71 $376.66\pm 1.16$
fastvit_sa36 41,065,117 256 29.9528 1,051.02 $44.58\pm 1.6$
1024 479.1497 14,353.52 $421.28\pm 1.33$
fastvit_ma36 57,758,965 256 42.0794 1,282.94 $56.72\pm 1.84$
1024 673.1369 17,062.69 $676.72\pm 1.19$

### v2

@@ 4.1 Datasets
The PanNuke dataset [14] is the primary resource for training and evaluating our model. It comprises 189,744 annotated nuclei across 7,904 images, each of size 256$\times$256 pixels, spanning 19 distinct tissue types and categorized into five unique cell classes. These cell images were captured at a 40$\times$ magnification with a fine resolution of 0.25 $\mu$m/px. Notably, the dataset exhibits a significant class imbalance; particularly, the nuclei class of dead cells is markedly underrepresented, evident from the nuclei and tissue class statistics.
@@
The MoNuSeg dataset [15] is employed as a supplementary resource for nuclei segmentation. Unlike PanNuke, MoNuSeg is considerably smaller and does not categorize nuclei into various classes. In this study, only the test subset of MoNuSeg is used to assess our model. This subset includes 14 high-resolution images (1000 $\times$ 1000 px) captured at 40$\times$ magnification and a resolution of 0.25 $\mu$m/px, containing over 7,000 annotated nuclei spanning seven organ types (kidney, lung, colon, breast, bladder, prostate, and brain) across various disease states. Due to the absence of nuclei labels, classification performance cannot be evaluated with this dataset.
@@
The CoNSeP dataset [11], curated by Graham et al., comprises 41 H&E-stained colorectal adenocarcinoma whole slide images (WSIs) at a resolution of 0.25 $\mu$m/px, resized to 1024 $\times$ 1024 px to facilitate processing. This diverse dataset features stromal, epithelial, muscular, collagen, adipose, and tumorous regions. It also includes a variety of nuclei types derived from different originating cells, such as normal epithelial, dysplastic epithelial, inflammatory, necrotic, muscular, fibroblast, and miscellaneous nuclei types, including necrotic and mitotic cells, which aids in comprehensive phenotypic analysis.
@@
The GLySAC dataset [16], short for Gastric Lymphocyte Segmentation and Classification, focuses on segmenting and classifying nuclei within gastric pathology. It contains 59 H&E stained image tiles, each 1000x1000 pixels, sourced from gastric adenocarcinoma WSIs and captured at a 40$\times$ magnification using an Aperio digital scanner. The dataset encapsulates a total of 30,875 nuclei, categorized into three primary groups: Lymphocytes (12,081 nuclei), Epithelial nuclei (12,287 nuclei), encompassing both cancerous and normal cells, and Miscellaneous other nuclei types (6,507 nuclei).
@@ 4.2 Evaluation metrics
The Panoptic Quality (PQ) metric provides a comprehensive evaluation by combining two essential aspects given by Detection Quality (DQ) and Segmentation Quality (SQ). DQ reflects how well the model detects and correctly identifies individual nuclei, calculated as denoted in Equation 6, where $TP$, $FP$, and $FN$ represent the true positives, false positives, and false negatives, respectively.
@@
SQ assesses the accuracy of the segmentation for the detected nuclei, computed as the mean IoU (Intersection over Union) of matched pixels, as denoted in Equation 7, where $y$ and $\hat{y}$ denote the ground truth and predicted segments, respectively.
@@
Therefore, PQ is the product of detection and segmentation quality, as denoted in equation 8.
@@
Several metrics commonly utilized in machine learning were employed to evaluate instance classification performance. Precision ($P$), which quantifies the accuracy of the positive predictions, is defined in the Equation 9.
@@
Where $TP$ represents true positives and $FP$ represents false positives. Recall ($R$), also known as sensitivity, measures the ability of the model to detect all relevant instances, defined in Equation 10
@@
With $FN$ indicating false negatives. The F1 Score, a harmonic mean of precision and recall that balances these metrics is crucial in uneven class distribution, shown in Equation 11.
@@
Accuracy, indicating the overall correctness of the model, is formulated in Equation 12:
@@
To detail performance assessment in multi-class settings, the F1 Score is further refined through equations that include terms for each class $c$ and $d$, illustrating both traditional components and inter-class effects, shown in Equations 13, 14, and 15
@@
$F1_{c}=\frac{2(T_{Pc}+T_{Nc})}{2(T_{Pc}+T_{Nc})+2FP_{c}+2FN_{c}+FP_{d}+FN_{d}}$ (15)
@@ 4.3 Results on PanNuke
DIST 0.4390 0.2900 0.3430 0.2750 0.0000
Mask-RCNN 0.4720 0.4030 0.2900 0.3000 0.0690
Micro-Net 0.5040 0.4420 0.3330 0.3340 0.0510
HoVer-Net 0.5510 0.4910 0.4170 0.3880 0.1390
HoVer-UNet 0.5240 0.4780 0.4010 0.3790 0.0760
CellViT256 0.5670 0.5590 0.4050 0.4050 0.1440
CellViT-SAM-H 0.5810 0.5830 0.4170 0.4230 0.1490
NuLite-T 0.5722 0.5622 0.4155 0.4062 0.1370
NuLite-M 0.5752 0.5693 0.4308 0.4070 0.1379
@@
Tissue HoVer-Net CellViT256 CellViT-SAM-H NuLite-T NuLite-M NuLite-H
@@
No Reparameterized FastViT-T8 0.649 0.477 0.822 0.563 0.546 0.407 0.399 0.135 0.687 0.697 0.577 0.519 0.349
FastViT-T12 0.653 0.484 0.823 0.569 0.558 0.417 0.404 0.145 0.696 0.716 0.575 0.523 0.350
FastViT-S12 0.653 0.485 0.824 0.573 0.561 0.412 0.408 0.128 0.697 0.718 0.575 0.521 0.358
FastViT-SA12 0.654 0.482 0.824 0.568 0.562 0.412 0.403 0.142 0.696 0.715 0.581 0.525 0.366
FastViT-SA24 0.658 0.488 0.825 0.575 0.565 0.417 0.408 0.135 0.701 0.724 0.579 0.524 0.341
FastViT-SA36 0.660 0.490 0.827 0.574 0.575 0.417 0.410 0.125 0.704 0.730 0.582 0.523 0.367
FastViT-MA36 0.659 0.493 0.826 0.579 0.577 0.420 0.413 0.132 0.706 0.730 0.584 0.529 0.367
Reparameterized FastViT-T8 0.649 0.477 0.822 0.563 0.548 0.408 0.401 0.141 0.701 0.725 0.581 0.526 0.353
FastViT-T12 0.653 0.484 0.823 0.570 0.561 0.414 0.399 0.150 0.696 0.716 0.575 0.523 0.350
FastViT-S12 0.654 0.485 0.824 0.572 0.562 0.416 0.406 0.137 0.697 0.718 0.575 0.521 0.358
FastViT-SA12 0.654 0.482 0.824 0.567 0.554 0.417 0.398 0.145 0.696 0.715 0.581 0.525 0.366
FastViT-SA24 0.658 0.488 0.825 0.574 0.566 0.418 0.406 0.149 0.701 0.724 0.579 0.524 0.341
FastViT-SA36 0.660 0.490 0.827 0.575 0.569 0.431 0.407 0.138 0.704 0.730 0.582 0.523 0.367
FastViT-MA36 0.659 0.493 0.826 0.577 0.571 0.417 0.413 0.145 0.706 0.730 0.584 0.529 0.367
@@
We used the PanNuke dataset to evaluate the performance of our models on nuclei instance segmentation and classification of five distinct cell types: neoplastic, epithelial, inflammatory, connective, and dead cells. We consider the PQ and F-score for each nuclei type and the F-score for detection to perform a robust analysis. Furthermore, we compare our results with DIST, MASK-RCNN, MICRO-Net, HoVer-UNet, CellViT256, and CellViT-SAM-H. In the following results, we consider three NuLite versions: NuLite-T, NuLite-M, and NuLite-H, which respectively use FastViT-S12, FastViT-SA36, and FastViT-MA36 as encoders and, in inference, they are reparameterized as described in [13]. In the ablation study section, we justify the selected encoders and reparameterization. To compare our model with others, we take into account three aspects. First, we analyze the PQ for each nucleus type, reported in 2, then we analyze the F1-score, reported in Table 3. Lastly, we analyze the results regarding binary Panoptic Quality (bPQ) and multi-class Panoptic Quality (mPQ) over each tissue. All results are an average of over three training sessions, as the authors’ dataset suggested.
As notable, in Table 2, the CellViT-SAM-H model outperforms all models. However, our NuLite versions follow closely, particularly excelling in Inflammatory, where NuLite-M has the highest values with a PQ of 0.4373. Overall, NuLite models, particularly NuLite-H, show competitive results for PQ metrics that outperform all other models. In binary detection, Table 3, the CellViT-SAM-H model and the NuLite-M and H models achieve the highest F1-scores of 0.83, showing strong detection capabilities. The NuLite-H model performs exceptionally well for Neoplastic and Epithelial nuclei, with F1-scores of 0.71 and 0.73, respectively, aching the SOTA results. The CellViT256 and NuLite models show strong results in nucleus types like inflammatory, connective, and dead nuclei. Notably, NuLite-H achieves top F1 Scores in classifying Connective nuclei, while NuLite-M and NuLite-H excel in the Dead nuclei category. Despite the CellViT-SAM-H model being the best model performer, NuLite models, especially NuLite-H, perform almost equally. They often rank second and show strengths in specific areas like binary detection and classification of more challenging nuclei types. Lastly, we also compared our versions with CellViT over tissue types, as shown in Table 4. Again, these results demonstrate that NuLite in all its versions is similar to CellViT in binary panoptic quality (bPQ) and multi-class panoptic quality (mPQ). Therefore, these results show that using NuFast-ViT-H, we practically obtain the same results as CellViT-SAM-H in terms of PQ and F1-score. Also, considering a tiny version of NuLite, the results are not different from CellViT-SAM-H and are better or equal to CellViT256, which is the lightest version of it.
Table 6: Comparison between CellViT and NuLite (ours) over two input shapes (256, 1024), in terms of the number of parameters, number of multiplications and additions, and estimated size GPU latency
Model Encoder # Parameters (M) GLOPS Estimated Total Size (MB) GPU Latency (ms)
256 1024 256 1024 256 1024
CellViT ViT-256 46.75 132.89 2,125.94 1,859.98 26,953.06 $35.71\pm 0.37$ $1169.7\pm 148.92$
SAM-H 699.74 214.20 3,413.41 6,002.34 45,612.96 $103.89\pm 0.97$ $2389.14\pm 150.18$
NuLite FastViT-T8 5.28 10.83 173.22 380.01 5,764.12 $13.42\pm 0.77$ $178.89\pm 18.05$
FastViT-T12 10.13 19.36 309.70 528.54 7,850.22 $14.87\pm 0.45$ $214.77\pm 23.66$
FastViT-S12 12.05 19.76 316.16 546.18 8,017.28 $14.76\pm 0.41$ $212.3\pm 21.4$
FastViT-SA12 14.16 19.76 316.18 555.41 8,038.31 $14.78\pm 0.83$ $212.98\pm 23.6$
FastViT-SA24 24.13 21.46 343.22 715.31 9,999.14 $21.84\pm 0.35$ $267.83\pm 24.81$
FastViT-SA36 34.10 23.15 370.25 875.21 11,959.97 $29.99\pm 1.79$ $310.44\pm 24.64$
FastViT-MA36 47.93 32.54 520.45 1,067.91 14,214.10 $33.37\pm 1.34$ $446.3\pm 35.25$
NuLite-Rep FastViT-T8 5.26 10.82 173.17 341.02 5,141.21 $9.11\pm 0.54$ $159.97\pm 18.11$
FastViT-T12 10.09 19.35 309.65 472.35 6,952.55 $10\pm 0.27$ $187.04\pm 20.67$
FastViT-S12 12.01 19.76 316.11 489.99 7,119.61 $9.96\pm 0.23$ $189.04\pm 16.61$
FastViT-SA12 14.13 19.76 316.13 501.34 7,174.21 $10.45\pm 0.27$ $197.35\pm 19.55$
FastViT-SA24 24.08 21.45 343.16 623.45 8,531.03 $14.69\pm 0.86$ $225.49\pm 18.37$
FastViT-SA36 34.04 23.14 370.20 745.57 9,887.84 $18.66\pm 0.4$ $266.82\pm 19.13$
FastViT-MA36 47.85 32.53 520.39 913.95 11,753.44 $23.05\pm 0.86$ $402.67\pm 30.99$
@@ 4.4 Ablation study
In this section, we describe the methodology used for selecting our models. We evaluated our NuLite using all versions of FastViT, considering both reparameterized and non-reparameterized variants during the inference phase. As shown in Table 5, the results indicate that reparameterization does not significantly affect performance. Consequently, we opted for the reparameterized versions due to their superior computational efficiency, as demonstrated in Table 6. To provide a comprehensive representation, we selected three versions of NuLite, namely NuLite-T, NuLite-M, and NuLite-H, corresponding to tiny, medium, and large configurations, respectively. This selection was informed by the evaluation results on the PanNuke dataset, detailed in Table 5. The metrics used for evaluation included binary panoptic quality (PQ), multi-class panoptic quality (mPQ), F1-score, panoptic quality, and F1-score for each nucleus type in the PanNuke dataset. The nucleus types are denoted as Neoplastic (N), Epithelial (E), Inflammatory (I), Connective/Soft Tissue (C), and Dead (D). Upon analyzing the results for the reparameterized models, it is apparent that the performance metrics are closely aligned across most variants. Thus, model selection also considered inference time and model complexity, as detailed in Table 6. First, we excluded the results with FastViT-T8 because despite being the lightest, the results were the worst. Then, we grouped by GLOPS the rest of the models and obtained three groups, namely FastViT-T12, FastViT-S12, FastViT-SA12 as tiny models, FastViT-SA24 and FastViT-SA36 for medium models, and FastViT-MA36 as a huge model. Subsequently, we chose FastViT-S12 as the backbone for NuLite-T(iny), FastViT-SA24 as the backbone of NuList-M(medium), and FastViT-SA36 for NuLite-H(uge), each one for the best performance in its group.